OpenStudy (anonymous):

if 4f(x)+f(5-x)=x^2 what is f(x)?

OpenStudy (asnaseer):

this is almost identical to one of your previous problems where you were shown how to solve it: http://openstudy.com/users/zahra.qatar#/updates/4edaadb8e4b01b4ec4921a72

OpenStudy (dumbcow):

\[f(x) = \frac{1}{5}x^{2}+\frac{2}{3}x-\frac{5}{3}\]